Calculate the ^H for the dissolution reaction of salts below using the thermodynamic data given

a)LiCl -408.61kj/mol

b)NaBr -360.0 kJ/mol


1
Expert's answer
2021-04-13T08:04:08-0400

= LiCl-408.61/6 ×10-6

= -68.10167 × 10-6 J

= NaBr -360/6 × 10-6

= -60 × 10-6J